# Value Residual Learning For Alleviating Attention Concentration In Transformers

### v3

@@ Abstract
Transformers can capture long-range dependencies using self-attention, allowing tokens to attend to all others directly. However, stacking multiple attention layers leads to attention concentration. One natural way to address this issue is to use cross-layer attention, allowing information from earlier layers to be directly accessible to later layers. However, this approach is computationally expensive. To address this problem, we propose Transformer with residual value (ResFormer) which approximates cross-layer attention through adding a residual connection from the values of the the first layer to all subsequent layers. Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer. Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 10.4% fewer model parameters and 13.6% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer reduces KV cache size by nearly half with only a small performance penalty and can be integrated with other KV-efficient methods, yielding further reductions in KV cache, with performance influenced by sequence length and cumulative learning rate. Further visualization results suggest that Resformer and SVFormer alleviate attention concentration in deeper layers through avoiding value-state drains and enhance representation across most layers.  [^1] Code is available at https://github.com/Zcchill/Value-Residual-Learning.
Figure 1: (Left) The average entropy of token importance and the average hidden-state similarity for a randomly initialized 468M model. (Middle) The average entropy of token importance across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023). (Right) The average entropy of token importance across layers in ResFormer vs. the vanilla Transformer, where token importance is derived from the attention matrix. Lower entropy indicates more focused attention on specific tokens. More details can be found in Eqn. 11.
@@ 1 Introduction
The Transformer (Vaswani et al., 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin et al., 2019, Lan et al., 2020, Brown et al., 2020) and computer vision tasks (Dosovitskiy et al., 2021). The discovery of scaling laws (Hoffmann et al., 2022, Kaplan et al., 2020) has driven the pursuit of larger Transformer models by increasing network depth and width.
(a) Token importance.
(b) Norms of value states.
(c) Norms of hidden states.
Figure 2: The token importance (Xiao et al., 2024), value-state norms (Guo et al., 2024b), and hidden-state norms (Sun et al., 2024) of the first token across layers of 468M models. More Visualization results are available in Appendix A.2.
Training large models presents significant challenges. Balancing the depth and width of a Transformer model within a fixed parameter budget is particularly difficult. While research indicates that deeper models generalize more compositionally than shallower ones (Petty et al., 2024), the training and deployment of deep models remain problematic. Although Transformers use residual connections (He et al., 2016) to address the vanishing gradient issue, training very deep Transformers is still challenging. For example, a 32-layer Vision Transformer (ViT) may perform worse than a 24-layer one (Zhou et al., 2021). This is mainly due to the smoothing mechanism of attention (Shi et al., 2022), which can lead to an over-smoothing effect (Nguyen et al., 2023) where the token representations become the same as the model’s depth increases.
Existing solutions to alleviate the over-smoothing problem in Transformer include adding extra regularizers (Nguyen et al., 2023, Shi et al., 2022) and optimizing the information flow within the model (Pagliardini et al., 2024). During the era of convolutional neural network architectures, Stochastic Depth (Huang et al., 2016) reduces the likelihood of over-smoothing by randomly dropping layers during training and DenseNet (Huang et al., 2017) mitigates the impact of over-smoothing by allowing each layer to directly access the hidden states of all preceding layers. Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer. Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value vectors of the first layer and the current layer to the attention output.
In this paper, we address the problem of multi-layer attention from another perspective. We introduce the phenomenon of attention concentration, which describes how a model’s attention increasingly focuses on fewer tokens. We quantify the degree of attention concentration using the entropy of the distribution of token importance, where lower entropy indicates a more pronounced concentration. Unlike over-smoothing, which is inherent to model architecture, attention concentration emerges during training. Fig. 1 (Left) shows that randomly initialized models exhibit over-smoothing but not attention concentration. Trained ViT models often focus on low-informative background areas (Darcet et al., 2024), while language models concentrate on low-semantic tokens (Sun et al., 2024), particularly the start token (attention sink (Xiao et al., 2024)). While previous studies analyzed single-layer attention patterns, our research reveals a “concentration - dispersion - concentration” pattern in deep models, as shown in Fig. 1 (Middle), suggesting potential loss of information during concentrated phases. The analysis of over-smoothing is available in Appendix A.1.
Mitigating attention concentration can lead to more interpretable attention maps and potentially improve downstream task performance (Darcet et al., 2024). This phenomenon typically emerges after the second or third network layer and is associated with value-state drains (decreased magnitude of value states) (Guo et al., 2024b), and hidden-state peaks (increased magnitude of hidden states) (Sun et al., 2024). Guo et al. (2024a) shows a mutual reinforcement mechanism exists between value-state drains and attention concentration. Recent studies have linked this to implicit biases during pretraining, with most existing solutions focusing on the use of additional tokens (registers) (Darcet et al., 2024) or additional keys and values (explicit attention bias) (Sun et al., 2024) to redirect this.
Given that the first layer always shows no attention concentration, an effective method is to use cross-layer attention on information from this layer. However, due to computational costs, we propose ResFormer as an efficient alternative. ResFormer applies a residual connection between the value vectors of the current layer and the first layer before the attention operation. Unlike cross-layer attention, ResFormer indirectly mitigates attention concentration. It leverages the absence of value-state drains in the first layer by introducing a value residual connection. This alleviates value-state drains in deeper layers, thereby disrupting the mutual reinforcement between attention concentration and value-state drains, as shown in Fig. 1 (Right) and Fig. 2.
During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2024). Existing $KV$-efficient methods often process keys and values simultaneously. Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer with single layer value (SVFormer). In SVFormer, the queries and keys of all layers share the value from the first layer, and thus it can also alleviate attention concentration.
We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023, Dubey et al., 2024, Kaplan et al., 2020). We compare different models by their relative training curves against the vanilla Transformer. Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO. ResFormer achieves equivalent validation loss with 10.4% fewer model parameters and 13.6% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer. And the performance of SVFormer is better when the training sequence length is longer. It further reduces the $KV$ cache when integrated with classical method GQA (Ainslie et al., 2023).
@@
Figure 3: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention. $A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively. $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
@@ 2 Related Work
2.1 Shortcut Connections For Better Information Flow
@@
Related research indicates that for advanced Transformer architectures, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024). Furthermore, Zhou et al. (2021) illustrates that a 32-layer ViT underperforms a 24-layer ViT. Depth-Wise Attention (ElNokrashy et al., 2024) allows each query to access the key and value at the same position from previous layers through an attention-like mechanism before the output layer. DenseFormer (Pagliardini et al., 2024) integrates weighted fusion of outputs from all preceding layers after each layer. To explore why increasing depth in Transformers does not yield expected gains, Wang et al. (2022) finds that self-attention acts as a low-pass filter, smoothing token representations in ViTs. Additionally, Shi et al. (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks. NeuTRENO (Nguyen et al., 2023) adds the difference between the value vectors of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
In contrast to these methods, ResFormer accesses and integrates information from previous layers prior to the attention operation, as illustrated in Fig. 3. Moreover, it does not require the selection or tuning of additional hyperparameters.
@@ 2.2 $KV$ cache compressing
The $KV$ cache is a key factor limiting the efficiency of long-text model inference. Research in this area can be broadly classified into Transformer-based methods, which target redundant information in Transformer models, and non-Transformer methods, which mainly addresses the quadratic time complexity of attention with respect to sequence length.
For non-Transformer methods, Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023) are two popular works. They replace the original softmax-based attention with SSM (Gu et al., 2021) and AFT (Zhai et al., 2021) mechanisms, respectively. Besides, several approaches have been proposed to enhance models’ ability to process long texts while reducing the reliance on $KV$ cache. Dai et al. (2019) advocates segmenting long texts into smaller parts for attention computation. Furthermore, Munkhdalai et al. (2024) uses a fixed-size memory matrix for storing and retrieving past information.
Transformer-based methods can be categorized into three main groups. The first group consists of post-training methods like SnapKV (Li et al., 2024) and ThinK (Xu et al., 2024), which compress $KV$ cache during inference based on attention matrices at token or hidden dimension levels. The second group focuses on quantization and adopts low-precision $KV$ cache quantization rather than completely eliminating them (Hooper et al., 2024). The third group aims to maximize the efficiency of attention-based models via parameter or activation value sharing. The most representative works include Multi-Query Attention (Shazeer, 2019) and Grouped-Query Attention (Ainslie et al., 2023) which suggest to share key and value across a group of queries. MLKV (Zuhri et al., 2024) further suggest to share keys and values for queries across layers and MLA (Liu et al., 2024) introduces low-rank projection when processing keys and values. Besides, CLA (Brandon et al., 2024) and LISA (Mu et al., 2024) respectively point out that we can reuse keys, values, or the attention matrix across layers to reduce redundancy between layers. While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key during attention computation. Moreover, it is compatible with other methods like GQA.
@@ 3 Method
3.1 Motivation: Information Transfer Via Cross Layer Attention
@@
Let $\mathbf{H}_{n}\in\mathbb{R}^{l\times d}$ be the input hidden state of the $n$-th layer, where $l$ denotes the sequence length and $d$ is the dimension size. In standard attention, the hidden state $\mathbf{H}_{n}$ will be firstly projected into $\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}\in\mathbb{R}^{l\times d}$ through three linear projections $\mathbf{W^{\mathbf{Q}}},\mathbf{W^{\mathbf{K}}},\mathbf{W^{\mathbf{V}}}\in\mathbb{R}^{d\times d}$ respectively. For simplicity, we introduce dot-product attention of layer $n$ as
$\operatorname{Attention}(\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n})$ $=\operatorname{Softmax}(\frac{\mathbf{Q}_{n}\mathbf{K}_{n}^{T}}{\sqrt{d}})\mathbf{V}_{n}.$ (1)
An ideal way to incorporate previous layers’ information is cross layer attention. The attention mechanism naturally extracts relevant information from previous layers. If these layers contain low-quality information, the similarity between the current layer’s query and the previous layers’ keys will be low, thus minimizing negative impacts. Given $m<n$ and the information $(\mathbf{Q}_{m},\mathbf{K}_{m},\mathbf{V}_{m})$ of $m$-th layer, the cross layer mechanism calculates the attention output $\mathbf{U}_{n}$ of $n$-th layer by the following attention formula:
$\mathbf{U}_{n}$ $=\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\operatorname{Concat}(\mathbf{K}_{n},\mathbf{K}_{m})^{T}}/{\sqrt{d}}\Big{)}\operatorname{Concat}(\mathbf{V}_{n},\mathbf{V}_{m}).$ (2)
In practice, cross-layer attention enhances feature fusion by allowing information to flow between layers, capturing both intra-layer and inter-layer dependencies. However, this approach introduces additional computational overhead due to the concatenation of keys and values from multiple layers. For example, in scenarios described by Eqn. 2, the overall computational complexity of the model nearly doubles compared with vanilla attention described in Eqn. 1.
3.2 Efficient Cross Layer Attention
To solve this problem, we propose to replace the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$ in Eqn. 2, as shown in Eqn. 3.
$\mathbf{U}_{n}$ $\approx\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\operatorname{Concat}(\mathbf{K}_{n},\mathbf{K}_{n})^{T}}/{\sqrt{d}}\Big{)}\operatorname{Concat}(\mathbf{V}_{n},\mathbf{V}_{m})$ (3)
$=\frac{1}{2}\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\mathbf{K}_{n}^{T}}/{\sqrt{d}}\Big{)}({\mathbf{V}_{n}}+{\mathbf{V}_{m}}).$ (4)
Utilizing the concept of block matrices, Eqn. 3 can be further simplified into Eqn. 4. This simplification converts the concatenation operation of the two value matrices into an addition operation. Compared to Eqn. 1, this new method only brings a minimal increase in computational complexity while still leveraging the information from the $m$-th layer in the $n$-th layer. Furthermore, Eqn. 4 can be generalized to incorporate cross-layer attention across all preceding $n-1$ layers as follows:
$\mathbf{U}_{n}$ $\approx\frac{1}{n}\mathbf{A}_{n}\sum_{i=1}^{n}{\mathbf{V}_{i}}.$ (5)
where $\mathbf{A}_{n}$ denotes the original attention matrix for layer $n$. From the perspective of information propagation, model described by Eqn. 3 projects the historical values into the current layer’s embedding space using the current layer’s attention as a weight matrix. For example, a naive approach would be to perform identity mapping, as described by $\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6) To evaluate the approximation effect of replacing the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$, we randomly select 1,000 pre-training data samples. For each layer of a trained baseline model, assuming cross-layer attention is required for each layer with respect to the previous one, we calculate the cosine similarity between the outputs from Eqn. 2 and Eqn. 4. We also calculate the cosine similarity between the outputs from Eqn. 2 and Eqn. 6 for comparison. Fig. 4 shows that our proposed method provides a good approximation for cross-layer attention.     [Uncaptioned image] Figure 4: Average token similarity between layer outputs. Lines show similarity of outputs using current attention (Eqn. 4) or identity attention (Eqn. 6) compared to the one using cross-layer attention (Eqn. 2).
3.3 Transformer with Residual Value
Based on Eqn. 5, we propose a variant of Transformer with residual value (ResFormer) which only chooses first layer as the target of cross layer attention since the first layer contains all basic information of each token. The analysis of entropy in Fig. 1 (Right) supports this point, indicating that attention tends to be relatively dispersed across different tokens in the initial layers of the model. The attention mechanism of ResFormer can be formulated as
$\mathbf{U}_{n}$ $=\frac{1}{2}\mathbf{A}_{n}({\mathbf{V}_{n}}+{\mathbf{V}_{1}}).$ (7)
where $n\geq 2$ and standard attention is applied in the first layer. From the training perspective, it explicitly learns a residual mapping instead of directly learning the desired underlying mapping and that’s why we call it ResFormer.
3.4 A unified View Of NeuTRENO and DenseFormer
Using our framework, the NeuTRENO can be defined as
$\mathbf{U}_{n}$ $=\Big{(}\mathbf{A}_{n}-\bm{\lambda}\mathbf{I}\Big{)}{\mathbf{V}_{n}}+\bm{\lambda}{\mathbf{V}_{1}}.$ (8)
where $\mathbf{I}$ denotes the identity matrix and $\bm{\lambda}$ is a hyper-parameter. It can be found that the term of $\bm{\lambda}\mathbf{I}$ may have certain negative impact on the learning of original attention. If we ignore the attention output projection and the MLP layer, DenseFormer can also be modeled within our framework as
$\mathbf{U}_{n}$ $=\sum_{i=1}^{n}{\bm{\alpha}_{i}\mathbf{A}_{i}\mathbf{V}_{i}}.$ (9)
where $\{\bm{\alpha}_{i}\}_{i=1}^{n}$ is a set of hyper-parameters. DenseFormer uses attention matrix of previous layer as the weight matrix of projecting values but this is not aligned with the concept shown in Eqn. 3.
3.5 SVFormer: Single-layer Value For Half $KV$ Cache
After ResFormer, a natural idea is whether we can remove the value vectors in each layer and have all layers share the value vectors from the first layer. We call this method SVFormer. Similar to ResFormer, SVFormer still adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through
$\mathbf{U}_{n}$ $=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ (10)

### v4

@@ Abstract
While Transformer models have achieved remarkable success in various domains, the effectiveness of information propagation through deep networks remains a critical challenge. Standard hidden state residuals often fail to adequately preserve initial token-level information in deeper layers. This paper introduces ResFormer, a novel architecture that enhances information flow by incorporating value residual connections in addition to hidden state residuals. And a variant is the SVFormer, where all layers share the first layer’s value embedding. Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 13.3% fewer model parameters and 15.4% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer reduces KV cache size by nearly half with only a small performance penalty and can be integrated with other KV-efficient methods, yielding further reductions in KV cache, with performance influenced by sequence length and cumulative learning rate.
Value Residual Learning
Zhanchao Zhou1,2   Tianyi Wu3†thanks: Equal Contribution; †Work done during internship at Westlake University; $\diamond$ Corresponding author.   Zhiyun Jiang4†∗   Fares Obeid5   Zhenzhong Lan2⋄ 1Zhejiang University      2Westlake University 3University of Electronic Science and Technology of China 4China University of Mining and Technology     5Imperial College London
@@ 1 Introduction
The Transformer (Vaswani et al., 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin et al., 2019; Lan et al., 2020; Brown et al., 2020) and computer vision tasks (Dosovitskiy et al., 2021). Among its variants, decoder-only architectures have become the most prominent (Kaplan et al., 2020; Dubey et al., 2024). The discovery of scaling laws (Hoffmann et al., 2022; Kaplan et al., 2020) has driven the pursuit of larger Transformer models by increasing network depth and width.
In a standard decoder-only transformer, initial token embeddings contain localized information, which rapidly e volves into abstract semantic features through early attention layers (Sun et al., 2024b; Clark et al., 2019). As Transformer deepens, a critical question arises: How effectively is the initial information propagated to deeper layers? One common answer is that residual connections of hidden states ensure access to initial information throughout the network. However, some studies (Zhou et al., 2021; Shi et al., 2022) have identified that the smoothing effect of attention mechanisms leads to over-smoothing, where token representations become increasingly similar as the network deepens. This indicates that in deeper layers, sequence-level features become dominant, while token-level features are diluted. DenseFormer (Pagliardini et al., 2024) applied the idea of learnable dense connections from DenseNet (Huang et al., 2016) to transformers, and the learned connection coefficients shows that deeper layers indeed require larger attention to initial embeddings. Given the low similarity between initial token embeddings and deeper hidden states (Sun et al., 2024b), their directly summation may significantly impact the modeling of attention distribution for abstract semantic information in later layers. NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing from the view of regularizers by considering the difference between value vectors of the first and current layers.
@@
Figure 1: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention. $\mathbf{A}^{i}$, $\mathbf{V}^{i}$, and $\mathbf{H}^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively. $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
In this paper, we propose ResFormer, which enhances the propagation of initial local information by introducing value residual connections in addition to the standard hidden residual connections. Specifically, ResFormer applies a residual connection between the value vectors of the current layer and the first layer before the attention operation. In other words, both value states share the existing attention matrix of the current layer. The value states of the first attention layer and the preceding hidden states differ only by a linear transformation along the channel dimension, both representing token-level raw information. We hypothesize that introducing residual connections for values has a less impact on modeling attention distributions for sequence-level semantic information in higher layers and complements the original hidden state residual. Fig. 1 illustrates a comparison of the extra skip connections introduced by different models.
During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2024). Existing $KV$-efficient methods often process keys and values simultaneously. Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer (SVFormer) where all layers share a single value states.
We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023; Dubey et al., 2024; Kaplan et al., 2020). We compare different models based on the valid loss against the vanilla Transformer. Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO. ResFormer achieves equivalent validation loss with 13.3% fewer model parameters and 15.4% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer. And SVFormer performs better when the training sequence length is longer. It further reduces the $KV$ cache when integrated with GQA (Ainslie et al., 2023).
@@ 2 Related Work
2.1 Shortcut Connections
@@
Related research indicates that, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024). Furthermore, Zhou et al. (2021) illustrates that a 32-layer ViT underperforms a 24-layer ViT. DenseFormer (Pagliardini et al., 2024) integrates weighted fusion of outputs from all preceding layers after each layer. To explore why increasing depth in Transformers does not yield expected gains, Wang et al. (2022) finds that self-attention acts as a low-pass filter, smoothing token representations in ViTs. Additionally, Shi et al. (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks. NeuTRENO (Nguyen et al., 2023) adds the difference between the value vectors of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
@@ 2.2 $KV$ cache compressing
The KV cache significantly impacts the efficiency of long-text model inference, attracting extensive research. One category of Transformer-based methods addresses this by employing parameter or activation value sharing techniques. The most representative works include Multi-Query Attention (Shazeer, 2019) and Grouped-Query Attention (Ainslie et al., 2023) which suggest to share key and value across a group of queries. Besides, CLA (Brandon et al., 2024) and LISA (Mu et al., 2024) respectively point out that we can reuse keys, values, or the attention matrix across layers to reduce redundancy between layers. While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key during attention.
@@ 3 Method
3.1 Notations
@@
Let $\mathbf{H}_{n}\in\mathbb{R}^{l\times d}$ be the input hidden state of the $n$-th layer, where $l$ denotes the sequence length and $d$ is the dimension size. For each layer, the hidden state $\mathbf{H}_{n-1}$ will be firstly projected into $\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}\in\mathbb{R}^{l\times d}$ through three linear projections $\mathbf{W^{\mathbf{Q}}_{n}},\mathbf{W^{\mathbf{K}}_{n}},\mathbf{W^{\mathbf{V}}_{n}}\in\mathbb{R}^{d\times d}$ respectively. After these projections, the attention operation ($\operatorname{Attn}$), output projection ($\mathbf{W}^{\mathbf{O}}_{n}\in\mathbb{R}^{d\times d}$), and Multi-Layer-Perceptron ($\operatorname{Mlp}$) are applied sequentially:
$\mathbf{U}_{n}$ $=\operatorname{Attn}(\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}).$ (1)
$\mathbf{H}_{n}$ $=\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n}).$ (2)
3.2 NeuTRENO and DenseFormer
After Eqn. 1, NeuTRENO adds the difference between the first and current layer’s value:
$\mathbf{U}_{n}$ $=\operatorname{Attn}(\mathbf{Q}_{n},\mathbf{K}_{n},\!\mathbf{V}_{n})+{\bm{\lambda}_{n}}(\mathbf{V}_{1}-\mathbf{V}_{n}).$ (3)
After Eqn. 2, DenseFormer performs a weighted average between all previous hidden states:
$\mathbf{H}_{n}$ $=\bm{\lambda}_{n,n}\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n})+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{H}_{i}.$ (4)
where $\mathbf{H}_{0}=\operatorname{Embedding}(\mathbf{X})$ for the input $\mathbf{X}$. $\bm{\lambda}_{n}$ in Eqn. 3 and $\!\{\bm{\lambda}_{n}\!\}$ in Eqn. 4 are new parameters.
3.3 ResFormer
In contrast, before Eqn. 1, ResFormer introduces a skip connection from the first layer’s value $\mathbf{V}_{0}=\mathbf{H}_{0}\mathbf{W}^{\mathbf{V}}_{1}$ to current layer’s value $\mathbf{V}_{n}=\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}$:
$\mathbf{V}_{n}$ $={\bm{\lambda}_{n,1}}\mathbf{V}_{1}+{\bm{\lambda}_{n,2}}\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}.$ (5)
where ${\bm{\lambda}_{n,1}}$ and ${\bm{\lambda}_{n,2}}$ are flexible scalars.
When all $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are predetermined constants, it is termed Constant-ResFormer. If $\{{\bm{\lambda}_{n,1}}={\bm{\lambda}_{n,2}}\}_{n=1}^{N}$, where $N$ is the total number of layers, the model is called Identity-ResFormer. Another variant where some layers have $\bm{\lambda}_{1}=0$ is referred to as Sparse-ResFormer. Besides, if $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are trainable parameters, the model is termed Learnable-ResFormer. Unless otherwise specified, $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are initialized to 0.5 for Learnable-ResFormer and are predetermined as 0.5 for Identity-ResFormer.
A more general form is the Dense-ResFormer, defined as $\mathbf{V}_{n}\!=\!\bm{\lambda}_{n,n}\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{V}_{i}$ for $n\geq 2$, where $\{\bm{\lambda}_{n}\!\}$ are constants or trainable scalars. Unless noted, all $\bm{\lambda}$ are initialized to 1 here.
3.4 SVFormer
Shared Parts - values keys keys & values
Valid Loss 2.739 2.743 2.753 2.776
Table 1: Results of sharing different parts every 2 layers.